# Improving Transferable Targeted Attacks with Feature Tuning Mixup

### v1

@@ Abstract
Deep neural networks exhibit vulnerability to adversarial examples that can transfer across different models. A particularly challenging problem is developing transferable targeted attacks that can mislead models into predicting specific target classes. While various methods have been proposed to enhance attack transferability, they often incur substantial computational costs while yielding limited improvements. Recent clean feature mixup methods use random clean features to perturb the feature space but lack optimization for disrupting adversarial examples, overlooking the advantages of attack-specific perturbations. In this paper, we propose Feature Tuning Mixup (FTM), a novel method that enhances targeted attack transferability by combining both random and optimized noises in the feature space. FTM introduces learnable feature perturbations and employs an efficient stochastic update strategy for optimization. These learnable perturbations facilitate the generation of more robust adversarial examples with improved transferability. We further demonstrate that attack performance can be enhanced through an ensemble of multiple FTM-perturbed surrogate models. Extensive experiments on the ImageNet-compatible dataset across various models demonstrate that our method achieves significant improvements over state-of-the-art methods while maintaining low computational cost.
@@ 1 Introduction
In this work, we propose Feature Tuning Mixup (FTM), a novel attack method that optimizes attack-specific feature perturbations to enhance the transferability of targeted adversarial examples. Our method fine-tunes the intermediate features of surrogate models during the attack process by introducing learnable feature perturbations. These feature perturbations are optimized by mixing them with clean features from random images to counteract the adversarial effects of targeted adversarial examples. Through this mechanism, adversarial examples are compelled to gradually develop more robust perturbations that can overcome these feature-level disturbances. To address the computational overhead of optimizing feature perturbations, we propose a stochastic update strategy. This strategy initializes the learnable perturbations using values from previous attack iterations and selectively optimizes only a small subset of feature perturbations in each iteration. Ultimately, FTM significantly improves the transferability of targeted adversarial examples while maintaining computational efficiency.
We provide an overview of the proposed FTM in Figure 2. FTM first introduces learnable perturbations into the deep layers of the surrogate model. The method targets layers with small feature sizes to minimize computational overhead and avoid excessive perturbations that could hinder attack performance. The learnable perturbations are added element-wise to the original output features before passing to the next layer. In each attack iteration, a small subset of layers is randomly selected and combined with random clean features. Finally, we compute gradients for the learnable perturbations and adversarial image using only one forward and backward pass, similar to standard attacks. We update the selected learnable perturbations and adversarial image, while unselected perturbations remain unchanged. As a result, FTM fully leverages optimized attack-specific noises, prevents overfitting to surrogate models, and significantly enhances the transferability of targeted attacks.
@@
  • We conduct extensive experiments across various source and target models on the ImageNet-compatible dataset, demonstrating that our FTM achieves a significant improvement over state-of-the-art methods.
@@ 2 Related Work
Adversarial attacks [4, 12, 19, 27, 7, 26] have been widely studied in the field of adversarial machine learning. Notably, black-box attacks can be categorized into query-based [1] and transfer-based attacks [13, 30, 9]. The latter, which does not require access to the target model’s outputs, parameters, or architecture, is more practical and is the focus of this paper.
@@ 2.2 Targeted Transfer-based Attacks
Targeted attacks are more challenging than non-targeted attacks because they not only aim to cause incorrect predictions by the models but also to mislead the models into predicting a specific target class. We can categorize targeted attacks into two types based on whether they require a training dataset. Previous works have utilized training datasets to generate auxiliary modules [21, 22] for transfer-based attacks or directly train a model to generate adversarial examples [31, 45]. This paper primarily focuses on transfer-based attacks that do not require a training dataset.
@@ 3.1 Preliminaries
$x^{\text{adv}}_{i+1}=\text{Clip}_{x,\epsilon}\left(x^{\text{adv}}_{i+1}\right)$ (5)
@@
CFM. To improve the transferability of targeted attacks, the Clean Feature Mixup (CFM) method [3] leverages stored clean features to augment adversarial examples, achieving state-of-the-art performance among existing targeted transfer-based attacks. Specifically, it selects outputs from certain network layers as perturbation targets and uses output features from different benign images as perturbation noises. To integrate with MI-FGSM, the feature augmentation in CFM can be formulated as follows:
@@ 3.3 Feature Tuning Mixup (FTM)
7:     Augment input $x^{\text{adv}}_{i}=\phi(x^{\text{adv}}_{i})$
@@ 4.1 Experimental Settings
Source Attack VGG-16 RN-18 RN-50 DN-121 Xcep MB-v2 EF-B0 IR-v2 Inc-v3 Inc-v4 Avg.
RN-50 DI 63.2 56.4 98.9 74.7 5.5 27.8 29.0 5.6 9.0 9.4 37.9
RDI 66.3 70.7 98.2 81.9 13.2 46.4 47.2 17.1 29.9 23.3 49.4
RDI-SI 70.5 79.8 98.8 88.9 29.5 56.2 66.2 37.9 56.4 43.6 62.8
RDI-VT 68.8 78.7 98.2 82.5 27.9 54.5 56.1 32.8 45.8 37.9 58.3
RDI-Admix 74.2 80.7 98.7 86.8 20.9 59.4 56.1 26.7 42.7 34.1 58.0
RDI-SU 67.9 74.7 98.3 83.3 24.9 56.3 50.2 20.9 32.1 26.2 53.5
ODI 78.3 77.1 97.6 87.0 43.8 67.3 70.0 49.5 65.9 55.4 69.2
RDI-CFM 84.7 88.4 98.4 90.3 51.1 81.5 78.8 48.0 65.5 59.3 74.6
RDI-FTM 86.3 87.5 97.9 89.7 56.1 83.3 81.2 54.7 70.8 66.6 77.4
RDI-FTM-E 88.1 88.6 98.3 91.4 59.4 85.4 84.3 56.9 73.4 69.3 79.5
Inc-v3 DI 2.9 2.4 3.4 5.0 1.9 1.8 3.7 3.0 99.2 4.2 12.8
RDI 3.5 3.8 4.0 7.0 3.1 3.0 5.9 6.3 98.7 7.1 14.2
RDI-SI 4.0 5.2 5.7 11.0 6.3 4.6 8.2 11.6 98.8 12.1 16.8
RDI-VT 5.9 8.9 9.4 13.2 7.4 5.9 9.8 12.3 98.7 14.7 18.6
RDI-Admix 6.3 6.5 8.8 12.8 6.0 6.1 10.9 12.2 98.7 13.6 18.2
RDI-SU 7.8 11.0 11.4 12.7 9.4 8.2 6.3 8.3 98.7 10.9 18.5
ODI 14.3 14.9 16.7 32.3 20.3 13.7 25.3 26.4 95.9 31.6 29.1
RDI-CFM 22.9 26.8 26.2 39.1 34.1 27.1 38.6 36.2 95.9 44.8 39.2
RDI-FTM 25.8 30.9 32.1 41.7 36.3 30.7 41.4 42.9 94.1 48.1 42.4
RDI-FTM-E 34.2 39.3 40.6 54.4 47.0 37.1 51.6 52.4 96.6 59.2 51.2
Table 1: Targeted attack success rates (%) against ten target models on the ImageNet-Compatible dataset. All methods are combined with MI-TI. The best results are shown in bold, and the second best results are underlined.
@@
Settings for our FTM method. We set the mixing ratio $\beta$ to 0.01. For FTM-E, we set the number of perturbed models to 2 for all experiments due to the computational cost. We follow CFM to only perturb network layers where the feature map size is smaller than or equal to $\frac{1}{16}$ of the image size. We sample $\bm{\alpha}_{k,i}$ from $\mathcal{U}(0,0.75)$, i.e., $\alpha_{max}=0.75$. We set the probability $p$ to $0.1$.
@@ 4.2 Experimental Results
Source Attack ViT LeViT ConViT Twins PiT Avg.
Computation time
per image (sec)
RN-50 DI 0.2 3.5 0.4 1.5 1.7 1.5 1.37
RDI 0.7 13.2 1.7 6.1 7.0 5.7 1.23
RDI-SI 2.9 29.4 6.3 15.5 17.9 14.4 5.95
RDI-VT 2.9 28.1 5.2 15.0 14.0 13.0 7.16
RDI-Admix 1.3 22.5 2.5 8.5 8.4 8.6 3.55
RDI-SU 0.8 16.9 2.3 6.9 7.8 6.9 2.21
ODI 5.1 37.0 10.7 20.1 29.1 20.4 4.38
RDI-CFM 4.3 46.1 8.9 25.2 24.7 21.8 1.39
RDI-FTM 5.9 52.9 10.8 32.4 31.5 26.7 1.54
RDI-FTM-E 6.8 58.6 13.6 35.2 34.9 29.8 2.92
Inc-v3 DI 0.1 0.3 0.0 0.0 0.1 0.1 2.01
RDI 0.2 1.8 0.2 0.4 0.7 0.7 1.76
RDI-SI 0.3 4.1 0.9 0.7 3.2 1.8 8.11
RDI-VT 0.4 5.2 0.8 1.6 1.8 2.0 10.5
RDI-Admix 0.1 4.1 0.6 1.4 1.4 1.5 4.92
RDI-SU 0.2 2.0 0.3 1.3 1.0 1.0 2.36
ODI 0.8 12.4 1.7 3.5 6.7 5.0 6.42
RDI-CFM 2.1 21.9 3.2 6.1 11.6 8.9 2.13
RDI-FTM 2.4 25.0 4.5 10.0 15.3 11.5 2.35
RDI-FTM-E 3.8 32.7 6.8 12.7 20.4 15.3 4.37
Table 2: Targeted attack success rates (%) against five transformer-based classifiers with the ImageNet-Compatible dataset. We also report the average computation time to generate an adversarial example. All methods are combined with MI-TI. The best results are shown in bold, and the second best results are underlined.
@@
Attack Inc-v3 ViT LeViT ConViT Twins
ODI / +CFM / +FTM-E 55.4 / 60.7 / 69.8 5.1 / 8.6 / 9.9 37.0 / 48.7 / 59.2 10.7 / 16.7 / 16.9 20.1 / 30.1 / 37.5
RDI-VT / +CFM / +FTM-E 37.9 / 62.0 / 70.5 2.9 / 11.2 / 13.5 28.1 / 53.0 / 59.0 5.2 / 15.7 / 18.3 15.0 / 33.6 / 38.2
RDI-Admix / +CFM / +FTM-E 34.1 / 64.8 / 68.7 1.3 / 6.3 / 7.2 22.5 / 52.3 / 59.1 2.5 / 10.7 / 14.6 8.5 / 28.1 / 36.0
RDI-SI / +CFM / +FTM-E 43.6 / 69.7 / 73.9 2.9 / 12.4 / 14.1 29.4 / 60.1 / 64.5 6.3 / 16.7 / 19.5 15.5 / 39.7 / 43.8
Table 3: Targeted attack success rates (%) of combination with other methods. The surrogate model is RN-50. All methods are combined with MI-TI. The best results are shown in bold.
@@
Combination with existing techniques. Table 3 shows the transfer success rates of our method combined with existing methods, including ODI, VT, SI, and Admix. For instance, RDI-FTM-E achieves an average success rate of 36.0% on Twins when combined with Admix, which is 7.9% higher than the success rate of RDI-Admix-CFM alone (28.1%). The results indicate that our method is a generic technique that can be combined with existing attack methods. More experiments are provided in the supplementary material.
@@ 4.3 Ablation Study
Figure 6: Left: Attack success rates vs. $p$ and $\alpha_{max}$. Right: Attack success rates vs. $p$ with/without clean features.

### v2

@@ Abstract
Deep neural networks (DNNs) exhibit vulnerability to adversarial examples that can transfer across different DNN models. A particularly challenging problem is developing transferable targeted attacks that can mislead DNN models into predicting specific target classes. While various methods have been proposed to enhance attack transferability, they often incur substantial computational costs while yielding limited improvements. Recent clean feature mixup methods use random clean features to perturb the feature space but lack optimization for disrupting adversarial examples, overlooking the advantages of attack-specific perturbations. In this paper, we propose Feature Tuning Mixup (FTM), a novel method that enhances targeted attack transferability by combining both random and optimized noises in the feature space. FTM introduces learnable feature perturbations and employs an efficient stochastic update strategy for optimization. These learnable perturbations facilitate the generation of more robust adversarial examples with improved transferability. We further demonstrate that attack performance can be enhanced through an ensemble of multiple FTM-perturbed surrogate models. Extensive experiments on the ImageNet-compatible dataset across various DNN models demonstrate that our method achieves significant improvements over state-of-the-art methods while maintaining low computational cost.  [^1] Our code is available at https://github.com/uhiu/feature-tuning-mixup
@@ 1 Introduction
In this work, we propose Feature Tuning Mixup (FTM), a novel attack method that optimizes attack-specific feature perturbations to enhance the transferability of targeted adversarial examples. Our method fine-tunes the intermediate features of surrogate models during the attack process by introducing learnable feature perturbations. These feature perturbations are optimized by mixing them with clean features from random images to counteract the adversarial effects of targeted adversarial examples. Through this mechanism, adversarial examples are gradually compelled to become more robust, allowing them to overcome these feature-level disturbances. To address the computational overhead of optimizing feature perturbations, we propose a stochastic update strategy. This strategy initializes the learnable perturbations using values from previous attack iterations and selectively optimizes only a small subset of feature perturbations in each iteration. Ultimately, FTM significantly improves the transferability of targeted adversarial examples while maintaining computational efficiency.
We provide an overview of the proposed FTM in Figure 2. FTM first introduces learnable perturbations into the deep layers of the surrogate model. The learnable perturbations are added element-wise to the original output features before passing to the next layer. In each attack iteration, a small subset of layers is randomly selected and combined with random clean features. Finally, we compute gradients for the learnable perturbations and adversarial image using only one forward and backward pass, similar to standard attacks. We update the selected learnable perturbations and adversarial image, while unselected perturbations remain unchanged. As a result, FTM fully leverages optimized attack-specific feature perturbations, prevents overfitting to surrogate models, and significantly enhances the transferability of targeted attacks.
@@
  • We conduct extensive experiments across various source and target DNN models on the ImageNet-compatible dataset, demonstrating that our FTM achieves a significant improvement over state-of-the-art methods.
@@ 2 Related Work
Adversarial attacks [4, 12, 19, 13, 31] have been widely studied in the field of adversarial learning, including unrestricted attacks [7], physical attacks [26, 27], and black-box attacks [1, 28, 9]. Notably, black-box attacks can be categorized into query-based [1] and transfer-based attacks [9, 28]. The latter, which does not require access to the target DNN model’s outputs, parameters or architecture, is more practical and is the focus of this paper.
@@ 2.2 Targeted Transfer-based Attacks
Targeted attacks are more challenging than non-targeted attacks because they not only aim to cause incorrect predictions by the models but also to mislead the models into predicting a specific target class. We can categorize targeted attacks into two types based on whether they require a training dataset. Previous works have utilized training datasets to generate auxiliary modules [21, 22] for transfer-based attacks or directly train a model to generate adversarial examples [32, 47, 42]. This paper primarily focuses on transfer-based attacks that do not require a training dataset.
@@ 3.1 Preliminaries
$x^{\text{adv}}_{i+1}=\text{Clip}_{x,\epsilon}\left(x^{\text{adv}}_{i+1}\right),$ (5)
@@
CFM. To improve the transferability of targeted attacks, the Clean Feature Mixup (CFM) method [3] leverages stored clean features to augment adversarial examples, achieving state-of-the-art performance among existing targeted transfer-based attacks. Specifically, it selects outputs from certain network layers as perturbation targets and uses output features from different benign images as feature perturbations. To integrate with MI-FGSM, the feature augmentation in CFM can be formulated as follows:
@@ 3.3 Feature Tuning Mixup (FTM)
Efficiency analysis. FTM does not require complex operations or additional forward and backward passes. Therefore, FTM is theoretically efficient on both CNNs and ViTs. In contrast, SI, VT, Admix, and SU require additional forward and backward passes, while ODI requires a computationally expensive input transformation.
@@
7:     Extend $F(x^{\text{adv}}_{i})$ to $F(\phi(x^{\text{adv}}_{i}))$
@@ 4 Experiments
Source Attack Time (s) VGG-16 RN-18 RN-50 DN-121 Xcep MB-v2 EF-B0 IR-v2 Inc-v3 Inc-v4 Avg.
RN-50 DI 1.37 63.2 56.4 98.9 74.7 5.5 27.8 29.0 5.6 9.0 9.4 37.9
RDI 1.23 66.3 70.7 98.2 81.9 13.2 46.4 47.2 17.1 29.9 23.3 49.4
RDI-SI 5.95 70.5 79.8 98.8 88.9 29.5 56.2 66.2 37.9 56.4 43.6 62.8
RDI-VT 7.16 68.8 78.7 98.2 82.5 27.9 54.5 56.1 32.8 45.8 37.9 58.3
RDI-Admix 3.55 74.2 80.7 98.7 86.8 20.9 59.4 56.1 26.7 42.7 34.1 58.0
RDI-SU 2.21 67.9 74.7 98.3 83.3 24.9 56.3 50.2 20.9 32.1 26.2 53.5
ODI 4.38 78.3 77.1 97.6 87.0 43.8 67.3 70.0 49.5 65.9 55.4 69.2
RDI-CFM 1.39 84.7 88.4 98.4 90.3 51.1 81.5 78.8 48.0 65.5 59.3 74.6
RDI-FTM 1.54 86.3 87.5 97.9 89.7 56.1 83.3 81.2 54.7 70.8 66.6 77.4
RDI-FTM-E 2.92 88.1 88.6 98.3 91.4 59.4 85.4 84.3 56.9 73.4 69.3 79.5
Inc-v3 DI 2.01 2.9 2.4 3.4 5.0 1.9 1.8 3.7 3.0 99.2 4.2 12.8
RDI 1.76 3.5 3.8 4.0 7.0 3.1 3.0 5.9 6.3 98.7 7.1 14.2
RDI-SI 8.11 4.0 5.2 5.7 11.0 6.3 4.6 8.2 11.6 98.8 12.1 16.8
RDI-VT 10.50 5.9 8.9 9.4 13.2 7.4 5.9 9.8 12.3 98.7 14.7 18.6
RDI-Admix 4.92 6.3 6.5 8.8 12.8 6.0 6.1 10.9 12.2 98.7 13.6 18.2
RDI-SU 2.36 7.8 11.0 11.4 12.7 9.4 8.2 6.3 8.3 98.7 10.9 18.5
ODI 6.42 14.3 14.9 16.7 32.3 20.3 13.7 25.3 26.4 95.9 31.6 29.1
RDI-CFM 2.13 22.9 26.8 26.2 39.1 34.1 27.1 38.6 36.2 95.9 44.8 39.2
RDI-FTM 2.35 25.8 30.9 32.1 41.7 36.3 30.7 41.4 42.9 94.1 48.1 42.4
RDI-FTM-E 4.37 34.2 39.3 40.6 54.4 47.0 37.1 51.6 52.4 96.6 59.2 51.2
Table 1: Targeted attack success rates (%) against ten target models. All methods are combined with MI-TI. The best results are shown in bold, and the second best results are underlined. Time (s) denotes the average computation time required to attack a single image.
@@ 4.1 Experimental Settings
Settings for our FTM. We set the mixing ratio $\beta$ to 0.01. For FTM-E, we set the number of perturbed models to 2 for all experiments due to the computational cost. Following CFM, we avoid perturbing shallow layers because their low-level features vary significantly with input transforms and may cause excessive disturbance during attacks, as explained in CFM’s paper. Specifically, we perturb layers where the feature map size is smaller than or equal to $\frac{1}{16}$ of the image size. For a pure ViT, we select the outputs of all fully-connected layers. We sample $\bm{\alpha}_{k,i}$ from $\mathcal{U}(0,0.75)$, i.e., $\alpha_{max}=0.75$. We set the probability $p$ to $0.1$.
@@ 4.2 Experimental Results
Source Attack Time (s) ViT LeViT ConViT Twins PiT Avg.
RN-50 DI 1.37 0.2 3.5 0.4 1.5 1.7 1.5
RDI 1.23 0.7 13.2 1.7 6.1 7.0 5.7
RDI-SI 5.95 2.9 29.4 6.3 15.5 17.9 14.4
RDI-VT 7.16 2.9 28.1 5.2 15.0 14.0 13.0
RDI-Admix 3.55 1.3 22.5 2.5 8.5 8.4 8.6
RDI-SU 2.21 0.8 16.9 2.3 6.9 7.8 6.9
ODI 4.38 5.1 37.0 10.7 20.1 29.1 20.4
RDI-CFM 1.39 4.3 46.1 8.9 25.2 24.7 21.8
RDI-FTM 1.54 5.9 52.9 10.8 32.4 31.5 26.7
RDI-FTM-E 2.92 6.8 58.6 13.6 35.2 34.9 29.8
Inc-v3 DI 2.01 0.1 0.3 0.0 0.0 0.1 0.1
RDI 1.76 0.2 1.8 0.2 0.4 0.7 0.7
RDI-SI 8.11 0.3 4.1 0.9 0.7 3.2 1.8
RDI-VT 10.5 0.4 5.2 0.8 1.6 1.8 2.0
RDI-Admix 4.92 0.1 4.1 0.6 1.4 1.4 1.5
RDI-SU 2.36 0.2 2.0 0.3 1.3 1.0 1.0
ODI 6.42 0.8 12.4 1.7 3.5 6.7 5.0
RDI-CFM 2.13 2.1 21.9 3.2 6.1 11.6 8.9
RDI-FTM 2.35 2.4 25.0 4.5 10.0 15.3 11.5
RDI-FTM-E 4.37 3.8 32.7 6.8 12.7 20.4 15.3
Table 2: Targeted attack success rates (%) against five transformer-based DNNs. All methods are combined with MI-TI. The best results are shown in bold, and the second best are underlined. Time (s) denotes the average time required to attack a single image.
@@
Attack RN-50 + Inc-v3 $\Rightarrow$
Time (s) VGG-16 RN-50* Inc-v3* DN-121 IR-v2 Inc-v4 Xcep ViT LeViT ConViT Twins PiT Avg.
RDI 3.26 66.9 98.4 98.7 83.1 35.6 42.5 25.4 0.8 22.4 3.2 9.8 10.8 41.5
ODI 7.17 70.9 94.0 96.1 83.4 65.4 69.9 57.6 9.1 52.9 18.6 28.6 42.8 57.4
RDI-CFM 4.39 86.1 98.3 97.0 92.1 71.7 79.9 72.0 10.3 66.5 20.3 41.9 45.8 65.2
RDI-ODI-CFM 7.73 63.9 88.6 81.3 74.1 54.3 56.0 54.3 12.6 49.9 20.9 29.6 41.0 52.2
RDI-Admix-CFM 13.38 86.2 98.1 97.7 92.7 75.1 81.6 76.1 14.0 73.1 24.8 50.7 53.9 68.7
RDI-SI-CFM 22.13 87.5 98.1 97.3 92.7 78.0 81.6 78.0 20.8 76.3 32.9 55.0 59.3 71.5
RDI-FTM 4.47 87.5 97.8 96.7 92.4 75.4 81.7 74.0 14.5 71.0 25.2 49.3 54.4 68.3
RDI-FTM-E 8.81 88.2 98.1 97.1 93.4 76.7 83.4 77.1 18.4 74.9 30.1 53.4 58.0 70.7
RDI-Admix-FTM 14.10 86.0 96.7 96.5 92.2 78.3 83.8 77.3 20.1 75.8 29.9 55.7 58.5 70.9
RDI-SI-FTM 22.66 87.5 96.8 96.2 93.3 79.0 83.6 79.1 25.7 77.7 37.5 59.9 64.3 73.4
RDI-SI-FTM-E 43.43 87.8 97.3 96.7 93.9 79.1 83.5 78.2 27.3 79.2 42.9 62.9 65.6 74.5
Table 3: Targeted attack success rates (%) of combination with different methods. The surrogate model is the ensemble of RN-50 and Inc-v3. All methods are combined with MI-TI. The best results are shown in bold, and the second best are underlined. Time (s) denotes the average time required to attack a single image.
@@
Combination with existing techniques. Table 3 presents transfer success rates of our method combined with ensemble-based method [30], SI, and Admix. Using the ensemble of RN-50 and Inc-v3 as the surrogate model, our RDI-SI-FTM-E achieves 74.5% average success rate, surpassing existing methods. FTM notably enhances performance on transformer-based models. For instance, RDI-SI-FTM achieves 64.3% on PiT, which is 5.0% higher than RDI-SI-CFM (59.3%) with comparable computational cost. These results demonstrate that FTM is a generic technique that combines effectively with existing attack methods.
@@ 4.3 Ablation Study
Figure 6: Left: Attack success rates vs. $p$ and $\alpha_{max}$. Right: Attack success rates vs. $p$ with/without clean features; The blue line is the success rates when the clean features from CFM are integrated with our FTM, while the orange line represents only using our learnable perturbations.
@@
The drop of the blue line in Figure 6 occurs because the clean accuracy of the perturbed surrogate model becomes extremely low (3.8%) at $p=1.0$ when using clean feature mixup. Without clean feature mixup (orange line), the accuracy remains high at 65.3%. The low accuracy makes it impossible to generate effective adversarial examples, causing the attack success rate of the blue line to drop to near 0.